OpenStudy (anonymous):

You have a torn tendon and are facing arthroscopic surgery to fix it. The surgeon explains the risks of the surgery. Infection occurs in 2% of all cases and the repair fails in 11% of the cases. 0.5% of the time the repair fails and infection occurs. What is the probability that the operation is successful and infection-free?

jimthompson5910 (jim_thompson5910):

for something like this, I would draw out a probability tree 

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Here's how I got those numbers first there's a 0.5% chance that there's an infection and the repair fails, so if you get an infection, then there's a 0.5/2 = 0.25 = 25% chance the repair fails so there's a 75% chance it doesn't fail meaning that the probability of getting an infection and not having the repair fail is 0.02*0.75 = 0.015 = 1.5%

jimthompson5910 (jim_thompson5910):

it says that 11% of the cases result in repairs failing, so the remaining bit must be 11 - 0.5 = 10.5% finally, all the values in the very far column must add to 100%, so 0.5 + 1.5 + 10.5 + x = 100 12.5 + x = 100 x = 100 - 12.5 x = 87.5 and that explains how I got the answer of 87.5%

jimthompson5910 (jim_thompson5910):

hopefully that makes sense

OpenStudy (anonymous):

ohhh. yeah, it does. thanks so much!!

jimthompson5910 (jim_thompson5910):

you're welcome